OpenStudy (anonymous):

Can someone please help me with these question. I really don't know how to do them 1. f-1(4) when f(x) = 3x minus 4, all over 5 2. y + 10 = 2y + 1

OpenStudy (anonymous):

1) if you solve \(\dfrac{3x-4}{5}=4\) what do you get for x?

OpenStudy (anonymous):

you get 3x=16 right?

OpenStudy (anonymous):

I ask for x, not 3x

OpenStudy (anonymous):

x =??

OpenStudy (anonymous):

so x=16/3

OpenStudy (anonymous):

simplify more, it is x =8, ok?

OpenStudy (anonymous):

wait how does x=8?

OpenStudy (anonymous):

oh, I am sorry, not 8, you are correct, 16/3

OpenStudy (anonymous):

dont you have to do something with the -1 in f-1(4)?

OpenStudy (anonymous):

f^-1(4)

OpenStudy (anonymous):

oh no!! you confuse me, we don't get that number oh boy! \(\dfrac{3x-4}{5}=4\rightarrow 3x-4=20\rightarrow 3x = 24 \rightarrow x =8\) hahaha...

OpenStudy (anonymous):

ok??

OpenStudy (anonymous):

i thought u had to do something with the inverse of -1. or is that just there.

OpenStudy (anonymous):

yeah, be patient, please, now I do it

OpenStudy (anonymous):

\(f(\color{red}{x})=4\\f^{-1}(4) =\color{red}{x}\) you have \(f(\color{red}{8})=4\\f^{-1}(4) =\color{red}{8}\) right?

OpenStudy (anonymous):

oh i see

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

so the answer is 8 overal

OpenStudy (anonymous):

yup, and don't forget the way how to get the answer.

OpenStudy (anonymous):

i got it thanks :)

OpenStudy (anonymous):

:)